1. Rank in order, from largest to smallest, the potential energies Ua to Us of these four pairs of charges. Each + symbol represents the same amount of charge. 2r 2r
Answer #1

option 3 is correct .

Potential energy U = kq1q2/d

Ua = ke^2/r = u

Ub = ke*2e/r = 2u

Uc = k*e*2e/2r = u

Ud = k*(2e)^2/2r = 2u